# The Pólya-Tchebotarev problem with semiclassical external fields

### v1

@@ 4.1 Finding a max-min candidate
For every maximizing candidate $F_{0}$ and small enough $t$, $F_{0}^{t}$ is an element of $\mathcal{F}_{M}$. Because the energy of $F_{0}$ is maximal within $\mathcal{F}_{M}$, we obtain directly from Definition 3.12 and Theorem 3.13 that $D_{h}I^{\varphi}(F_{0})=0$, concluding the proof. ∎
@@ 4.2 Constructing a solution for the max-min problem
If $z\in F_{0}\setminus\operatorname{supp}\mu_{0}$ then $U^{\mu}+\varphi$ is harmonic near $z$, so it cannot have a local maximum in $z$. This implies that
@@
$S_{j}^{\epsilon}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\left\{z\in\mathbb{C};|z|\geq R,|\arg z-\theta_{j}|\leq\frac{\pi}{2N}+\epsilon\right\},\quad j=1,\ldots,N.$
For each $\epsilon>0$, there exists $R_{\epsilon}>0$ such that for every $R>R_{\epsilon}$ the set $\Lambda\setminus D_{R}$ has exactly $N$ connected components $\lambda_{1},\ldots,\lambda_{N}$, where
@@
Now, we fix $\epsilon>0$, also ensuring that the sets $S_{j}^{\epsilon}$ are pairwise disjoint. In addition, fix $R>R_{\epsilon}$ sufficiently large so that $\mathcal{Z}\cup\mathcal{C}\cup\operatorname{supp}\mu_{0}\subset D_{R}$. For $1\leq j\leq N$, let $\Lambda(j)$ be the connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ for which
@@
The set $\Lambda$ is open and has finitely many components whose boundary consists of a finite number of smooth arcs, whereas $\operatorname{supp}\mu_{0}$ consists of finitely many analytic arcs. These properties ensure that $\Lambda(j)$ is pathwise connected, and we can connect any point of $\Lambda(j)$ to infinity through a curve entirely contained in $\Lambda(j)$ and stretching out to infinity in the sector $S_{j}$. The following lemma characterizes $\Lambda(j)$ in terms of the fixed connections. With a slight abuse of notation, for a fixed point $c\in\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ we denote by $\Lambda(c)$ the connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ that contains $c$.
@@
From Lemma 4.4 we see that $x_{j}\in\Lambda(j)$. Let $C\in\mathcal{P}(\mathcal{C})$ with $\Psi(C)=\emptyset$. Then Lemma 4.5 implies that all points of $C$ are contained in the same connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. So there is a connected compact set $\gamma_{C}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$, which is a finite union of bounded piecewise smooth arcs, such that $C\subset\gamma_{C}$. For each $A\in\mathcal{P}(\mathcal{C})$ with $\Psi(A)\neq\emptyset$, Lemma 4.5 implies that all points of $A$ and $x_{j}$, with $\theta_{j}\in\Psi(A)$, are in the same connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. There is a connected compact set $\gamma_{A}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ which is a finite union of bounded piecewise smooth arcs, such that $A\subset\gamma_{A}$ and $x_{j}\in\gamma_{A}$ for each $\theta_{j}\in\Psi(A)$. For $L_{j,R}$ as in Lemma 4.4, define
@@
Each $\Gamma_{A}$ is a finite union of piecewise smooth arcs, $A\subset\Gamma_{A}$, $\Gamma_{A}$ is connected and stretches out to infinity on each sector $S_{j}$ with $\theta_{j}\in\Psi(A)$. Given $B\in\mathcal{P}(\Theta)\setminus\mathcal{P}_{\Psi}$, Lemma 4.5 implies that the points $x_{j}$ with $j\in B$ belong to the same connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. Then there exists a compact connected set $\gamma_{B}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$, which is a finite union of bounded piecewise smooth arcs, satisfying $x_{j}\in\gamma_{B}$ for each $\theta_{j}\in B$. Define
@@
Each $\Gamma_{B}$ is a finite union of piecewise smooth arcs. $\Gamma_{B}$ is connected and stretches out to infinity on each sector $S_{j}$ with $\theta_{j}\in B$. The union
$\Gamma_{0}=\left(\bigcup_{\begin{subarray}{c}C\in\mathcal{P}(\mathcal{C})\\ \Psi(C)=\emptyset\end{subarray}}\gamma_{A}\right)\bigcup\left(\bigcup_{\begin{subarray}{c}A\in\mathcal{P}(\mathcal{C})\\ \Psi(A)\neq\emptyset\end{subarray}}\Gamma_{A}\right)\bigcup\left(\bigcup_{B\in\mathcal{P}_{0}(\Theta_{\infty})}\Gamma_{B}\right)$
@@
and so equality holds throughout these inequalities. Let $\mu_{1}$ be the equilibrium measure of $\Gamma_{0}$ in the external field $\varphi$. Then
@@
Let $\Gamma_{n}$ be a sequence in $\mathcal{T}_{M}$ that converges to $F_{0}$ in the Hausdorff metric. Let $R$ be large enough such that the circle $|z|=R$ intersects each of the components of $\Delta_{M}$ along $N$ circular arcs. Then there are $N$ circular arcs of $|z|=R$ outside of $\Delta_{M}$, that we denote by $\eta_{1},\ldots,\eta_{N}$, where $\eta_{k}$ is in the direction of $S_{k}$. If $R$ is large enough, Lemma 4.4 ensures that
@@
for each $i$. We fix $R$ satisfying all conditions above and such that each fixed point $c$ is in $D_{R}$. We start by proving the second item. Let $c$ be a fixed point that $\mathcal{T}$ connects to infinity through $S_{j}$. For each $n\in\mathbb{N}$, $\Gamma_{n}$ has a connected component that connects $c$ to infinity through $S_{j}$. On this connected component we can find a subset $\Gamma_{n}^{\prime}$ which is a simple piece-wise $C^{1}$ contour connecting $c$ to $\infty$ through $S_{j}$. We consider $\Gamma_{n}^{\prime}$ as an oriented contour, going from $c$ to $\infty$, and say that $\alpha$ lies before a point $\omega\in\Gamma_{n}^{\prime}$ if $\alpha$ appears first when following the contour. By definition of $\mathcal{T}_{M}$,
@@
Now put $\eta=\cup_{i=1}^{N}\eta_{i}$. We follow the contour $\Gamma_{n}^{\prime}$ starting at $c$. Put $\omega_{0,n}=c$. Let $\alpha_{1,n}$ be the first point of $\Gamma_{n}^{\prime}$ after $\omega_{0,n}$ that intersects with $\eta$, then $\alpha_{1,n}\in\eta_{i_{1}}$ for some index $i_{1}$. If $i_{1}=j$ we stop our process. Otherwise, since $\Gamma_{n}^{\prime}$ goes to $\infty$ through $S_{j}$, there exists $\omega_{1,n}$ that is the last point of intersection of $\Gamma_{n}^{\prime}$ with $\alpha_{i_{1}}$. After that, $\Gamma_{n}^{\prime}$ enters $D_{R}$ and, since it must go to $\infty$ through $S_{j}$, it must intersect $\eta$ again. Let $\alpha_{2,n}$ be the first point after $\omega_{1,n}$ of this intersection, then $\alpha_{2,n}\in\alpha_{i_{2}}$ for some index $i_{2}\neq i_{1}$. If $i_{2}=j$ we are finished. Otherwise, there exists $\omega_{2,n}$ that is the last point of intersection of $\Gamma_{n}^{\prime}$ with $\alpha_{i_{2}}$. The process is clear, and can be followed until there is some $l\in\{1,\ldots,N-1\}$ where $i_{l}=j$. This produces indexes $i_{1},\ldots,i_{l}$, and points $c=\omega_{0,n},\alpha_{1,n},\omega_{1,n}\ldots,\omega_{l-1,n},\alpha_{l,n}$ on $\Gamma_{n}^{\prime}$ such that
@@
The numbers $i_{1},\ldots,i_{l-1}$ depends on the contour $\Gamma_{n}$, so they also depends on $n$. Since these numbers are integers between $1$ and $N$, we can pass to a subsequence and assume that they are independent of $n$. The set
$\Gamma_{n}^{\prime}[\omega_{k,n}\alpha_{k+1,n}]=\Gamma_{n}^{\prime}(\omega_{k,n},\alpha_{k+1,n})\cup\{\omega_{k,n},\alpha_{k+1,n}\}$
@@
since $\Gamma_{n}^{\prime}[\omega_{k,n},\alpha_{k+1,n}]\subset\Gamma_{n}$ for each $n$. By compactness of $\eta_{i}$ we can also assume (taking further subsequences) that $(\alpha_{i,n})_{n}$ and $(\omega_{i,n})_{n}$ converge to, say,
@@
which is exactly what we wanted to prove. The third case is similar. Let $c_{1},c_{2}$ be fixed points that are connected by $\mathcal{T}$. For each $n$ there exists a simple piece-wise $C^{1}$ contour $\gamma_{n}\subset\Gamma_{n}$ connecting $c_{1}$ to $c_{2}$. Assume that there exists a subsequence of $\gamma_{n}$ that does not intersect $\eta$. We pass to this subsequence, and thus $\gamma_{n}$ is contained in $D_{R}$ for each $n\in\mathbb{N}$. Going to a further subsequence we can assume that $\gamma_{n}$ converges to a set $\gamma$ in the Hausdorff metric. This $\gamma$ is a connected set that contains $c_{1}$ and $c_{2}$, and is contained in $F_{0}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. This obviously implies that $\Lambda(c_{1})=\Lambda(c_{2})$. If there is no subsequence as above then we can assume that $\gamma_{n}$ always intersect $\eta$. Then we do a similar construction as in the case of item $(2)$: Put $\omega_{0,n}=c_{1}$ and let $\alpha_{1,n}$ be the first point of intersection of $\gamma_{n}$ with $\eta$, then $\alpha_{1,n}\in\eta_{i_{1}}$ and, since $\gamma_{n}$ ends at $c_{2}$, there exists a last point of intersection $\omega_{1,n}$. If after $\omega_{1,n}$ the contour $\gamma_{n}$ does not intersect intersect $\eta$ anymore, we put $\alpha_{2,n}=c_{2}$ and we are done. Otherwise, let $\alpha_{2,n}$ be the first point of intersection of $\gamma_{n}$ with $\eta$, then $\alpha_{2,n}\in\eta_{i_{2}}$ and, since $\gamma_{n}$ ends at $c_{2}$, there exists a last point of intersection of $\gamma_{n}$ with $\eta_{i_{2}}$ that we denote by $\omega_{2,n}$. The process is clear, and we obtain a sequence $c_{1}=\omega_{0,n},\alpha_{1,n},\omega_{1,n},\ldots,\alpha_{i_{l},n},\omega_{i_{l},n},\alpha_{i_{l}+1,n}=c_{2}$. The rest of the proof follows by similar arguments as in the case of (2). Analogously we verify (1). ∎
@@ Acknowledgments
V. A. is supported by grant # 2020/13183-0, São Paulo Research Foundation (FAPESP) and grant # 2022/12756-1, São Paulo Research Foundation (FAPESP).

### v2

@@ 4.1 Finding a max-min candidate
For every maximizing candidate $F_{0}$ and small enough $t$, $F_{0}^{t}$ is an element of $\mathcal{F}_{M}$. Because the energy of $F_{0}$ is maximal within $\mathcal{F}_{M}$, we obtain directly from Definition 3.14 and Theorem 3.15 that $D_{h}I^{\varphi}(F_{0})=0$. Thus $F_{0}$ is a critical set, and the rest follows from Corollary 3.13, concluding the proof. ∎
@@ 4.2 Constructing a solution for the max-min problem
If $z\in F_{0}\setminus\operatorname{supp}\mu_{0}$ then $U^{\mu_{0}}+\varphi$ is harmonic near $z$, so it cannot have a local maximum in $z$. This implies that
@@
$S_{j}^{\epsilon}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\left\{z\in\mathbb{C};|z|\geq R,|\arg z-\theta_{j}|\leq\frac{\pi}{2N}+\epsilon\right\},\quad j=1,\ldots,N,$
where the angles $\theta_{j}$ were defined in (2.11). For each $\epsilon>0$, there exists $R_{\epsilon}>0$ such that for every $R>R_{\epsilon}$ the set $\Lambda\setminus D_{R}$ has exactly $N$ connected components $\lambda_{1},\ldots,\lambda_{N}$, where
@@
Now, we fix $\epsilon>0$, also ensuring that the sets $S_{j}^{\epsilon}$ are pairwise disjoint. In addition, fix $R>R_{\epsilon}$ sufficiently large so that $\mathcal{Z}\cup\mathcal{C}\cup\operatorname{supp}\mu_{0}\subset D_{R}$.
For $1\leq j\leq N$, let $\Lambda(j)$ be the connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ for which
@@
The set $\Lambda$ is open and has finitely many components whose boundary consists of a finite number of smooth arcs, whereas $\operatorname{supp}\mu_{0}$ consists of finitely many analytic arcs (see Lemma 4.3). These properties ensure that $\Lambda(j)$ is pathwise connected, and we can connect any point of $\Lambda(j)$ to infinity through a curve entirely contained in $\Lambda(j)$ and stretching out to infinity in the sector $S_{j}$.
The following lemma characterizes $\Lambda(j)$ in terms of the fixed connections. With a slight abuse of notation, for a fixed point $c\in\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ we denote by $\Lambda(c)$ the connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ that contains $c$.
@@
From Lemma 4.4 we see that $x_{j}\in\Lambda(j)$.
Let $C\in\mathcal{P}(\mathcal{C})$ with $\Psi(C)=\emptyset$. Then Lemma 4.5 implies that all points of $C$ are contained in the same connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. So there is a connected compact set $\gamma_{C}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$, which is a finite union of bounded piecewise smooth arcs, such that $C\subset\gamma_{C}$.
For each $A\in\mathcal{P}(\mathcal{C})$ with $\Psi(A)\neq\emptyset$, Lemma 4.5 implies that all points of $A$ and $x_{j}$, with $\theta_{j}\in\Psi(A)$, are in the same connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. There is a connected compact set $\gamma_{A}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$ which is a finite union of bounded piecewise smooth arcs, such that $A\subset\gamma_{A}$ and $x_{j}\in\gamma_{A}$ for each $\theta_{j}\in\Psi(A)$. For $L_{j,R}$ as in Lemma 4.4, define
@@
Each $\Gamma_{A}$ is a finite union of piecewise smooth arcs, $A\subset\Gamma_{A}$, $\Gamma_{A}$ is connected and stretches out to infinity on each sector $S_{j}$ with $\theta_{j}\in\Psi(A)$.
Given $B\in\mathcal{P}_{0}(\Theta)\setminus\mathcal{P}_{\Psi}$, Lemma 4.5 implies that the points $x_{j}$ with $j\in B$ belong to the same connected component of $\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. Then there exists a compact connected set $\gamma_{B}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$, which is a finite union of bounded piecewise smooth arcs, satisfying $x_{j}\in\gamma_{B}$ for each $\theta_{j}\in B$. Define
@@
Each $\Gamma_{B}$ is a finite union of piecewise smooth arcs. $\Gamma_{B}$ is connected and stretches out to infinity on each sector $S_{j}$ with $\theta_{j}\in B$.
The union
$\Gamma_{0}=\left(\bigcup_{\begin{subarray}{c}C\in\mathcal{P}(\mathcal{C})\\ \Psi(C)=\emptyset\end{subarray}}\gamma_{C}\right)\bigcup\left(\bigcup_{\begin{subarray}{c}A\in\mathcal{P}(\mathcal{C})\\ \Psi(A)\neq\emptyset\end{subarray}}\Gamma_{A}\right)\bigcup\left(\bigcup_{B\in\mathcal{P}_{0}(\Theta)\setminus\mathcal{P}_{\Psi}}\Gamma_{B}\right)$
@@
and so equality holds throughout these inequalities.
Let $\mu_{1}$ be the equilibrium measure of $\Gamma_{0}$ in the external field $\varphi$. Then
@@
Let $\Gamma_{n}$ be a sequence in $\mathcal{T}_{M}$ that converges to $F_{0}$ in the Hausdorff metric.
Let $R$ be large enough such that the circle $|z|=R$ intersects each of the components of $\Delta_{M}$ along $N$ circular arcs. Then there are $N$ circular arcs of $|z|=R$ outside of $\Delta_{M}$, that we denote by $\eta_{1},\ldots,\eta_{N}$, where $\eta_{k}$ is in the direction of $S_{k}$. If $R$ is large enough, Lemma 4.4 ensures that
@@
for each $i$. We fix $R$ satisfying all conditions above and such that each fixed point $c$ is in $D_{R}$.
We start by proving the second item. Let $c$ be a fixed point that $\mathcal{T}$ connects to infinity through $S_{j}$. For each $n\in\mathbb{N}$, $\Gamma_{n}$ has a connected component that connects $c$ to infinity through $S_{j}$. On this connected component we can find a subset $\Gamma_{n}^{\prime}$ which is a simple piece-wise $C^{1}$ contour connecting $c$ to $\infty$ through $S_{j}$. We consider $\Gamma_{n}^{\prime}$ as an oriented contour, going from $c$ to $\infty$, and say that $\alpha$ lies before a point $\omega\in\Gamma_{n}^{\prime}$ if $\alpha$ appears first when following the contour. By definition of $\mathcal{T}_{M}$,
@@
Now put $\eta=\cup_{i=1}^{N}\eta_{i}$. We follow the contour $\Gamma_{n}^{\prime}$ starting at $c$. Put $\omega_{0,n}=c$. Let $\alpha_{1,n}$ be the first point of $\Gamma_{n}^{\prime}$ after $\omega_{0,n}$ that intersects with $\eta$, then $\alpha_{1,n}\in\eta_{i_{1}}$ for some index $i_{1}$. If $i_{1}=j$ we stop our process. Otherwise, since $\Gamma_{n}^{\prime}$ goes to $\infty$ through $S_{j}$, there exists $\omega_{1,n}$ that is the last point of intersection of $\Gamma_{n}^{\prime}$ with $\alpha_{i_{1}}$. After that, $\Gamma_{n}^{\prime}$ enters $D_{R}$ and, since it must go to $\infty$ through $S_{j}$, it must intersect $\eta$ again. Let $\alpha_{2,n}$ be the first point after $\omega_{1,n}$ of this intersection, then $\alpha_{2,n}\in\alpha_{i_{2}}$ for some index $i_{2}\neq i_{1}$. If $i_{2}=j$ we are finished. Otherwise, there exists $\omega_{2,n}$ that is the last point of intersection of $\Gamma_{n}^{\prime}$ with $\alpha_{i_{2}}$.
The process is clear, and can be followed until there is some $l\in\{1,\ldots,N-1\}$ where $i_{l}=j$. This produces indexes $i_{1},\ldots,i_{l}$, and points $c=\omega_{0,n},\alpha_{1,n},\omega_{1,n}\ldots,\omega_{l-1,n},\alpha_{l,n}$ on $\Gamma_{n}^{\prime}$ such that
@@
The numbers $i_{1},\ldots,i_{l-1}$ depends on the contour $\Gamma_{n}$, so they also depends on $n$. Since these numbers are integers between $1$ and $N$, we can pass to a subsequence and assume that they are independent of $n$.
The set
$\Gamma_{n}^{\prime}[\omega_{k,n},\alpha_{k+1,n}]=\Gamma_{n}^{\prime}(\omega_{k,n},\alpha_{k+1,n})\cup\{\omega_{k,n},\alpha_{k+1,n}\}$
@@
since $\Gamma_{n}^{\prime}[\omega_{k,n},\alpha_{k+1,n}]\subset\Gamma_{n}$ for each $n$.
By compactness of $\eta_{i}$ we can also assume (taking further subsequences) that $(\alpha_{i,n})_{n}$ and $(\omega_{i,n})_{n}$ converge to, say,
@@
which is exactly what we wanted to prove.
The third case is similar. Let $c_{1},c_{2}$ be fixed points that are connected by $\mathcal{T}$. For each $n$ there exists a simple piece-wise $C^{1}$ contour $\gamma_{n}\subset\Gamma_{n}$ connecting $c_{1}$ to $c_{2}$.
Assume that there exists a subsequence of $\gamma_{n}$ that does not intersect $\eta$. We pass to this subsequence, and thus $\gamma_{n}$ is contained in $D_{R}$ for each $n\in\mathbb{N}$. Going to a further subsequence we can assume that $\gamma_{n}$ converges to a set $\gamma$ in the Hausdorff metric. This $\gamma$ is a connected set that contains $c_{1}$ and $c_{2}$, and is contained in $F_{0}\subset\overline{\Lambda}\cup\operatorname{supp}\mu_{0}$. This obviously implies that $\Lambda(c_{1})=\Lambda(c_{2})$.
If there is no subsequence as above then we can assume that $\gamma_{n}$ always intersect $\eta$. Then we do a similar construction as in the case of item $(2)$: Put $\omega_{0,n}=c_{1}$ and let $\alpha_{1,n}$ be the first point of intersection of $\gamma_{n}$ with $\eta$, then $\alpha_{1,n}\in\eta_{i_{1}}$ and, since $\gamma_{n}$ ends at $c_{2}$, there exists a last point of intersection $\omega_{1,n}$. If after $\omega_{1,n}$ the contour $\gamma_{n}$ does not intersect intersect $\eta$ anymore, we put $\alpha_{2,n}=c_{2}$ and we are done. Otherwise, let $\alpha_{2,n}$ be the first point of intersection of $\gamma_{n}$ with $\eta$, then $\alpha_{2,n}\in\eta_{i_{2}}$ and, since $\gamma_{n}$ ends at $c_{2}$, there exists a last point of intersection of $\gamma_{n}$ with $\eta_{i_{2}}$ that we denote by $\omega_{2,n}$. The process is clear, and we obtain a sequence $c_{1}=\omega_{0,n},\alpha_{1,n},\omega_{1,n},\ldots,\alpha_{i_{l},n},\omega_{i_{l},n},\alpha_{i_{l}+1,n}=c_{2}$. The rest of the proof follows by similar arguments as in the case of (2). Analogously we verify (1).
@@ Acknowledgments
Declarations.
Figures were made with Wolfram Mathematica 13.3 and Mathcha.io.
Ethical Approval. Not applicable.
Competing interests. The authors declare that they have no affiliations with or any type of involvement with any organization or entity with financial interest in the results discussed in this paper.
Author’s contributions. Victor Alves: Formal Analysis, Writing - original draft, Writing review and editing. Guilherme L.F. Silva: Formal Analysis, Writing - original draft, Writing review and editing.
Funding. V. A. is supported by grant # 2020/13183-0, São Paulo Research Foundation (FAPESP) and grant # 2022/12756-1, São Paulo Research Foundation (FAPESP).
@@
Availability of Data and materials. Data sharing is not applicable to this article since no datasets were created nor analysed.